# Capturing Polysemanticity with PRISM: A Multi-Concept Feature Description Framework

### v3

@@ 1 Introduction
A central goal of this research is to assign interpretable, functional roles to individual components such as neurons or attention heads bau-2017-NetworkDissectionQuantifying ; mu-2020-CompositionalExplanationsNeurons ; bykov-2023-LabelingNeuralRepresentations ; neo_interpreting_2024 . The presence of polysemanticity, the tendency of individual features to encode multiple, semantically distinct concepts or patterns scherlis-2022-PolysemanticityCapacityNeural ; elhage-2022-ToyModelsSuperposition , complicates the process of explaining model components, as it defies the common assumption that a single neuron is associated with a single function or pattern. From a coding-theoretic perspective, this reflects how neural capacity is distributed across multiple tasks. While several concept extraction techniques like sparse autoencoders (SAEs) ng2011sparse ; bricken-2023-MonosemanticityDecomposingLanguage aim to disentangle polysemantic features, many resulting descriptions still reflect multiple concepts openai2024automatedinterpretability , limiting interpretability.
While the problem of polysemanticity is generally addressed through the extraction of sparse features, such as via sparse autoencoders, feature description methods, which aim to explain the functional purpose of individual features, typically provide a single explanation per feature bills-2023-LanguageModelsCan ; gao-2024-ScalingEvaluatingSparse ; choi_scaling_2024 ; yu-2024-LatentConceptbasedExplanation ; gur-arieh_enhancing_2025 . This can limit the ability to capture the full range of patterns a feature may represent. To offer more comprehensive and nuanced feature descriptions, we introduce PRISM, a framework for generating and evaluating multi-concept feature descriptions that considers multiple patterns per feature. In Figure 1, we present a neuron previously labeled as “possibly monosemantic” openai2024automatedinterpretability that PRISM reveals to activate for a highly diverse and heterogeneous set of concepts characterized by textual descriptions. Throughout this work, a concept refers to a cluster of token-level inputs that elicit similar contextualized feature activations, allowing each feature to be associated with multiple concepts that reflect the diversity of inputs it responds to. Our contributions include:
@@ 2 Related Work
Recent generative AI systems have grown increasingly complex, motivating efforts to scale interpretability. In addition to methods that provide local explanations for individual predictions, a complementary line of research focuses on fully automated descriptions of model components like neurons and their associated functions. In this context, an automated interpretability approach was introduced to generate a textual description for each neuron in GPT-2 XL bills-2023-LanguageModelsCan . This method has since become a foundation for subsequent work on feature descriptions cunningham-2023-SparseAutoencodersFind ; bricken-2023-MonosemanticityDecomposingLanguage ; choi_scaling_2024 ; gao-2024-ScalingEvaluatingSparse ; he-2024-LlamaScopeExtracting ; gur-arieh_enhancing_2025 .
@@
Automated feature description methods often rely on simulation-based approaches, where a model predicts neuron activations for given inputs bills-2023-LanguageModelsCan . This approach, widely adopted in recent work cunningham-2023-SparseAutoencodersFind ; bricken-2023-MonosemanticityDecomposingLanguage ; choi_scaling_2024 , evaluates descriptions based on the correlation between simulated and actual activations. Other evaluations include binary classification of activating and non-activating contexts paulo-2024-AutomaticallyInterpretingMillions , contrastive methods using distractor samples mcgrath–MappingLatentSpace ; lindsey-2024-InterpretabilityEvalsDictionary , and combined approaches that assess input capture accuracy and feature characterization precision gur-arieh_enhancing_2025 ; huang-2023-RigorouslyAssessingNatural ; templeton-2024-scaling-monosemanticity . Simulation-based approaches herein rely on an external model’s ability to accurately predict neuron activations given a feature extracted by an additional explainer model. Currently, it remains unclear how well LLMs reliably predict activations of other models, adding uncertainty and complexity to the evaluation of feature descriptions. Our approach instead directly compares feature activation distributions to controls, using both parametric and non-parametric evaluation measures.
@@ 3.2 Evaluating Multi-Concept Feature Descriptions
Following previous work, which uses contrastive methods to differentiate between activating samples and control samples mcgrath–MappingLatentSpace ; lindsey-2024-InterpretabilityEvalsDictionary , we adapt the CoSy evaluation method kopf-2024-CoSyEvaluatingTextual to language models (see Figure 7 in Appendix A.3). CoSy evaluates each candidate feature using two complementary metrics. The Area Under the Receiver Operating Characteristic (AUC) measures how well the feature distinguishes between control data points $\mathbb{X}_{0}$, consisting of 1,000 randomly sampled entries from Cosmopedia benallalloubna-2024-Cosmopedia , and target concept data points $\mathbb{X}_{1}$, consisting of 10 concept-specific text samples generated by an LLM for the feature description. Given the corresponding activations $\mathbb{A}_{0}\in\mathbb{R}^{n}$ for $\mathbb{X}_{0}$ and $\mathbb{A}_{1}\in\mathbb{R}^{m}$ for $\mathbb{X}_{1}$, the AUC is computed as
$\Psi_{\text{AUC}}(\mathbb{A}_{0},\mathbb{A}_{1})=\frac{\sum_{a\in\mathbb{A}_{0}}\sum_{b\in\mathbb{A}_{1}}\mathbf{1}[a<b]}{|\mathbb{A}_{0}|\cdot|\mathbb{A}_{1}|}.$ (6)
@@
with mean control activation $\bar{a}=\frac{1}{n}\sum_{a\in\mathbb{A}_{0}}a$. See Appendix A.3 for dataset, model and activation details.
@@ 4.1 Experimental Setup
In our experiments, we evaluate PRISM against competitive feature description methods, GPT-Explain bills-2023-LanguageModelsCan and Output-Centric gur-arieh_enhancing_2025  [^5] We use descriptions generated by their Ensemble Raw (All) method, which is best performing on their input-based evaluation.. Unless otherwise specified, all experiments are conducted on the English training subset of the C4 corpus raffel-2020-ExploringLimitsTransfer , a large, cleaned version of Common Crawl’s web crawl corpus. We process text excerpts to a uniform length of 512 tokens, either by truncation or padding, ensuring consistency across model inputs. Feature activations are extracted from four pre-trained LLMs, from three layers per model (early, middle, late) [^6] More details can be found in GitHub Repository https://github.com/lkopf/prism. For GPT-2 XL radford-2019-LanguageModelsAre and Llama 3.1 8B Instruct grattafiori-2024-llama-3 , we analyze MLP neurons. For GPT-2 Small gao-2024-ScalingEvaluatingSparse and Gemma Scope lieberum-2024-GemmaScopeOpen , we focus on residual stream SAE features on account of the publicly available implementations. More details on the evaluation procedure and experimental setup are provided in Appendix A.3 and Appendix A.4.
@@
  • For (3) Cluster Labeling, to generate human-interpretable labels, we prompt a large language model (Gemini 1.5 Pro gemini-team-2024-gemini-1-5 ) using the $N_{s}=20$ text excerpts with the highest mean activations for each cluster. Only positive activations are considered, and a token is highlighted if its activation exceeds a sample-specific threshold, defined as the 90th percentile of the sample’s activation distribution. Token spans corresponding to peak activations are highlighted with square brackets “[…]” lee-2023-ImportancePromptTuninga ; paulo-2024-AutomaticallyInterpretingMillions ; choi_scaling_2024 in the prompt to guide the model. The LLM is instructed to output a concise summary of the shared concept in each cluster. Additional prompt details are provided in Appendix A.4.
@@ 4.3 Sanity Checks
Randomizing sentences or descriptions yields AUC values near random relative to the baseline (Table 6). While MAD scores are often not statistically significant due to large standard deviations, they consistently decrease, aligning with our expectations. For polysemanticity, randomly assigned descriptions yield notably lower similarity values than true scores, reflecting reduced semantic coherence (Figure 17(a)). Examining percentile intervals, the top 25% (0.75–1.0) of activations closely match baseline performance, whereas lower quartiles show reduced scores, demonstrating a positive correlation between activation strength and description quality (Table 7). The analysis of relative activations between the 99th and 100th percentiles shows that extremely small ratios are rare, especially in early layers, indicating that percentile-based sampling reliably captures diverse, meaningful patterns (Figure 20(a)).
@@ 4.4 Benchmarking Experiment Results
Table 1: Benchmarking feature description methods. AUC reflects classification performance, while MAD measures activation differences between target and control descriptions. PRISM (max) reports the best-matching score per feature; (mean) averages over all descriptions. Values are means across selected features from early, middle, and late layers. AUC includes 95% confidence intervals; MAD includes standard deviations. Bold indicates best performance; dashes denote unavailable descriptions for certain models. See Appendix A.4 for details.
@@
AUC ($\uparrow$) MAD ($\uparrow$) AUC ($\uparrow$) MAD ($\uparrow$) AUC ($\uparrow$) MAD ($\uparrow$) AUC ($\uparrow$) MAD ($\uparrow$)
[…]
@@
A key observation is that PRISM (max) achieves the highest AUC and MAD scores, across all model and feature-type combinations. For example, on GPT-2 XL (neuron features), PRISM (max) reaches an AUC of 0.85 and a MAD of 2.98, indicating that its descriptions are more accurate and outperform the competitive approach of GPT-Explain. One exception is observed with Gemma Scope (SAE features), where the Output-Centric method slightly outperforms PRISM (max), achieving the highest MAD (12.24) and AUC (0.58). This suggests that, for certain architectures or feature types, alignment in the output space may offer complementary advantages. Additional analysis on the distribution and variability of AUC and MAD scores is provided in Appendix A.4. Overall, we find that neuron-based features yield more reliable interpretations than SAE-based features, providing further evidence for recent methodological concerns about achieving improved interpretability through SAEs sharkey2025openproblemsmechanisticinterpretability .
@@
In Figure LABEL:fig:evaluation-scores and Figure LABEL:fig:poly-scores, we analyze the feature descriptions’ quality and polysemanticity across layers. Judging the evaluation scores, middle layers generally appear to be easier to interpret (Figure LABEL:fig:evaluation-scores). In most models, the AUC scores peak in the middle layer. An exception is observed in Llama 3.1 8B Instruct, where evaluation scores (measured in AUC, see 3.2) increase in the later layer.
(a) Comparison of PRISM (max) AUC evaluation scores and PRISM polysemanticity scores across different models and layers.
To analyze the degree of polysemanticity, we use the gte-Qwen2-1.5B-instruct sentence transformer li-2023-GeneralTextEmbeddings to embed the five natural language descriptions generated per feature. We then compute pairwise cosine similarity within each set of five descriptions to estimate semantic consistency. In Figure LABEL:fig:poly-scores, polysemanticity scores show no consistent trend across layers and vary across models and feature types. Notably, Gemma Scope SAE feature descriptions exhibit the highest monosemanticity across all layers. Interestingly, we can also note that Gemma Scope SAE feature descriptions show the highest monosemanticity across all layers, as shown in Figure LABEL:fig:poly-scores. Despite high variability, our findings suggest that polysemanticity does not consistently increase or decrease with layer depth, and interpretability varies significantly across architectures.
@@ 6 Discussion and Conclusion
As we have focused on textual explanations, our descriptions are limited to concepts expressible in natural language, and thus may be unable to capture complex syntactic structures like graphs or algorithmic concepts where the underlying operation is not easily described in the constrained vocabulary space. Fixing the number of clusters in PRISM cannot guarantee that provided descriptions capture all or the most salient concepts for a feature; instead, they represent a subset of key concepts the feature is responsive to. Parametric measures such as MAD are sensitive to outliers, especially in NLP settings where activations often follow heavy-tailed distributions. To address this, we pair MAD with the more robust AUC score, yielding a more comprehensive view of feature behavior. Finally, reliance on maximally activating corpus examples can restrict interpretations to observed concepts, limiting coverage of rare patterns or out-of-distribution effects across datasets bolukbasi_interpretability_2021 .

### v4

@@ 1 Introduction
A central goal of this research is to assign interpretable, functional roles to individual components such as neurons or attention heads bau-2017-NetworkDissectionQuantifying; mu-2020-CompositionalExplanationsNeurons; bykov-2023-LabelingNeuralRepresentations; neo_interpreting_2024. The presence of polysemanticity, the tendency of individual features to encode multiple, semantically distinct concepts or patterns scherlis-2022-PolysemanticityCapacityNeural; elhage-2022-ToyModelsSuperposition, complicates the process of explaining model components, as it defies the common assumption that a single neuron is associated with a single function or pattern. From a coding-theoretic perspective, this reflects how neural capacity is distributed across multiple tasks. While several concept extraction techniques like sparse autoencoders (SAEs) ng2011sparse; bricken-2023-MonosemanticityDecomposingLanguage aim to disentangle polysemantic features, many learned features still encode multiple concepts openai2024automatedinterpretability, and therefore cannot be regarded as truly monosemantic.
While the problem of polysemanticity is generally addressed through the extraction of sparse features, such as via sparse autoencoders, feature description methods, which aim to explain the functional purpose of individual features, typically provide a single explanation per feature bills-2023-LanguageModelsCan; gao_scaling_2024; choi_scaling_2024; yu-2024-LatentConceptbasedExplanation; gur-arieh_enhancing_2025. This can limit the ability to capture the full range of patterns a feature may represent. To offer more comprehensive and nuanced feature descriptions, we introduce PRISM, a framework for generating and evaluating multi-concept feature descriptions that considers multiple patterns per feature. In Figure 1, we present a neuron previously labeled as “possibly monosemantic” openai2024automatedinterpretability that PRISM reveals to activate for a highly diverse and heterogeneous set of concepts characterized by textual descriptions. Throughout this work, a concept refers to a cluster of token-level inputs that elicit similar contextualized feature activations, allowing each feature to be associated with multiple concepts that reflect the diversity of inputs it responds to. Our contributions include:
@@ 2 Related Work
Recent generative AI systems have grown increasingly complex, motivating efforts to scale interpretability. In addition to methods that provide local explanations for individual predictions, a complementary line of research focuses on fully automated descriptions of model components like neurons and their associated functions. In this context, an automated interpretability approach was introduced to generate a textual description for each neuron in GPT-2 XL bills-2023-LanguageModelsCan. This method has since become a foundation for subsequent work on feature descriptions cunningham-2023-SparseAutoencodersFind; bricken-2023-MonosemanticityDecomposingLanguage; choi_scaling_2024; gao_scaling_2024; he-2024-LlamaScopeExtracting; gur-arieh_enhancing_2025.
@@
Automated feature description methods often rely on simulation-based approaches, where a model predicts neuron activations for given inputs bills-2023-LanguageModelsCan. This approach, widely adopted in recent work cunningham-2023-SparseAutoencodersFind; bricken-2023-MonosemanticityDecomposingLanguage; choi_scaling_2024, evaluates descriptions based on the correlation between simulated and actual activations. Other evaluations include binary classification of activating and non-activating contexts paulo_automatically_2025, contrastive methods using distractor samples mcgrath--MappingLatentSpace; lindsey-2024-InterpretabilityEvalsDictionary, and combined approaches that assess input capture accuracy and feature characterization precision gur-arieh_enhancing_2025; huang-2023-RigorouslyAssessingNatural; templeton-2024-scaling-monosemanticity. Simulation-based approaches herein rely on an external model’s ability to accurately predict neuron activations given a feature extracted by an additional explainer model. Currently, it remains unclear how well LLMs reliably predict activations of other models, adding uncertainty and complexity to the evaluation of feature descriptions. Our approach instead directly compares feature activation distributions to controls, using both parametric and non-parametric evaluation measures.
@@ 3.2 Evaluating Multi-Concept Feature Descriptions
Following previous work, which uses contrastive methods to differentiate between activating samples and control samples mcgrath--MappingLatentSpace; lindsey-2024-InterpretabilityEvalsDictionary, we adapt the CoSy evaluation method kopf-2024-CoSyEvaluatingTextual to language models (see Figure 7 in Appendix A.3). CoSy evaluates each candidate feature using two complementary metrics. The Area Under the Receiver Operating Characteristic (AUROC) measures how well the feature distinguishes between control data points $\mathbb{X}_{0}$, consisting of 1,000 randomly sampled entries from Cosmopedia benallalloubna-2024-Cosmopedia, and target concept data points $\mathbb{X}_{1}$, consisting of 10 concept-specific text samples generated by an LLM for the feature description. Given the corresponding activations $\mathbb{A}_{0}\in\mathbb{R}^{n}$ for $\mathbb{X}_{0}$ and $\mathbb{A}_{1}\in\mathbb{R}^{m}$ for $\mathbb{X}_{1}$, the AUROC is computed as
$\Psi_{\text{AUROC}}(\mathbb{A}_{0},\mathbb{A}_{1})=\frac{\sum_{a\in\mathbb{A}_{0}}\sum_{b\in\mathbb{A}_{1}}\mathbf{1}[a<b]}{|\mathbb{A}_{0}|\cdot|\mathbb{A}_{1}|}.$ (6)
@@
with mean control activation $\bar{a}=\frac{1}{n}\sum_{a\in\mathbb{A}_{0}}a$. In our evaluation, we report the percentage of features with positive MAD scores, i.e., the fraction of features satisfying $\mathbb{A}_{1}>\mathbb{A}_{0}$. See Appendix A.3 for dataset, model and activation details.
@@ 4.1 Experimental Setup
In our experiments, we evaluate PRISM against competitive feature description methods, MaxAct [^5] For obtaining feature descriptions with the MaxAct method, we use a subset of 10,000 samples from the English training split of the C4 corpus raffel-2020-ExploringLimitsTransfer, collect the top five activating samples per feature, and generate feature descriptions using the same prompt as in PRISM(see Appendix A.4)., GPT-Explain bills-2023-LanguageModelsCan, Transluce-Explain choi_scaling_2024, Neuronpedia neuronpedia, and Output-Centric gur-arieh_enhancing_2025666We use descriptions generated by their Ensemble Raw (All) method, which is best performing on their input-based evaluation.. Unless otherwise specified, all experiments are conducted on the English training subset of the C4 corpus raffel-2020-ExploringLimitsTransfer, a large, cleaned version of Common Crawl’s web crawl corpus. We process text excerpts to a uniform length of 512 tokens, either by truncation or padding, ensuring consistency across model inputs. Feature activations are extracted from four pre-trained LLMs, from three layers per model (early, middle, late) [^7] More details can be found in GitHub Repository https://github.com/lkopf/prism. For GPT-2 XL radford-2019-LanguageModelsAre and Llama 3.1 8B Instruct grattafiori-2024-llama-3, we analyze MLP neurons. For GPT-2 Small gao_scaling_2024 and Gemma Scope lieberum-2024-GemmaScopeOpen, we focus on residual stream SAE features on account of the publicly available implementations. More details on the evaluation procedure and experimental setup are provided in Appendix A.3 and Appendix A.4.
@@
  • For (3) Cluster Labeling, to generate human-interpretable labels, we prompt a large language model (Gemini 1.5 Pro gemini-team-2024-gemini-1-5888The model originally employed in this study (Gemini 1.5 Pro) was deprecated upon completion of the project. Consequently, subsequent experiments, including those pertaining to MaxAct and output-centric evaluation, were conducted using the available and comparable model, Gemini 2.0 Flash Lite comanici2025gemini.) using the $N_{s}=20$ text excerpts with the highest mean activations for each cluster. Only positive activations are considered, and a token is highlighted if its activation exceeds a sample-specific threshold, defined as the 90th percentile of the sample’s activation distribution. Token spans corresponding to peak activations are highlighted with square brackets “[…]” lee-2023-ImportancePromptTuninga; paulo_automatically_2025; choi_scaling_2024 in the prompt to guide the model. The LLM is instructed to output a concise summary of the shared concept in each cluster. Additional prompt details are provided in Appendix A.4.
@@ 4.3 Sanity Checks
Randomizing sentences or descriptions yields AUROC values near random relative to the baseline (Table 6). While MAD scores are often not statistically significant due to large standard deviations, they consistently decrease, aligning with our expectations. For polysemanticity, randomly assigned descriptions yield notably lower similarity values than true scores, reflecting reduced semantic coherence (Figure 17(a)). Examining percentile intervals, the top 25% (0.75–1.0) of activations closely match baseline performance, whereas lower quartiles show reduced scores, demonstrating a positive correlation between activation strength and description quality (Table 7). The analysis of relative activations between the 99th and 100th percentiles shows that extremely small ratios are rare, especially in early layers, indicating that percentile-based sampling reliably captures diverse, meaningful patterns (Figure 20(a)).
@@ 4.4 Benchmarking Experiment Results
Table 1: Benchmarking feature description methods. AUROC reflects classification performance, while MAD measures activation differences between target and control descriptions. PRISM (max) reports the best-matching score per feature; (mean) averages over all descriptions. Values are means across selected features from early, middle, and late layers. AUROC includes 95% confidence intervals; MAD represents the percentage of positive MAD scores. Bold indicates best performance; dashes denote unavailable descriptions for certain models. See Appendix A.4 for details.
@@
AUROC ($\uparrow$) MAD ($\uparrow$) AUROC ($\uparrow$) MAD ($\uparrow$) AUROC ($\uparrow$) MAD ($\uparrow$) AUROC ($\uparrow$) MAD ($\uparrow$)
MaxAct 0.53 (0.49-0.58) 11.86% 0.54 (0.46-0.63) 50.00% 0.53 (0.49-0.58) 11.86% 0.60 (0.50-0.69) 50.00%
GPT-Explain bills-2023-LanguageModelsCan 0.64 (0.56-0.73) 65.00%
Transluce-Explain choi_scaling_2024 0.59 (0.51-0.67) 63.33%
Neuronpedia neuronpedia 0.54 (0.50-0.59) 18.97% 0.62 (0.53-0.72) 63.33%
Output-Centric gur-arieh_enhancing_2025 0.55 (0.46-0.64) 58.33% 0.57 (0.53-0.62) 22.03% 0.58 (0.49-0.67) 46.67%
PRISM (mean) 0.65 (0.61-0.69) 66.33% 0.52 (0.48-0.55) 51.33% 0.51 (0.50-0.53) 13.22% 0.43 (0.39-0.46) 24.67%
PRISM (max) 0.85 (0.78-0.91) 91.67% 0.71 (0.63-0.78) 81.67% 0.57 (0.53-0.61) 28.81% 0.54 (0.45-0.62) 38.33%
@@
A key observation is that PRISM (max) achieves the highest AUROC and MAD scores, across all model and feature-type combinations. For example, on GPT-2 XL (neuron features), PRISM (max) reaches an AUROC of 0.85 and a MAD of 91.67%, indicating that its descriptions are more accurate and outperform the competitive approach of GPT-Explain. One exception is observed with Gemma Scope (SAE features), where the Neuronpedia method slightly outperforms PRISM (max), achieving the highest MAD (63.33%) and AUROC (0.62). This suggests that, for certain architectures or feature types, alignment in the output space may offer complementary advantages. Additional analysis on the distribution and variability of AUROC and MAD scores is provided in Appendix A.4. Overall, we find that neuron-based features yield more reliable interpretations than SAE-based features, providing further evidence for recent methodological concerns about achieving improved interpretability through SAEs sharkey2025openproblemsmechanisticinterpretability. Extended results on the output-centric evaluation are provided in Appendix A.4.
@@
In Figure LABEL:fig:evaluation-scores and Figure LABEL:fig:poly-scores, we analyze the feature descriptions’ quality and polysemanticity across layers. Judging the evaluation scores, middle layers generally appear to be easier to interpret (Figure LABEL:fig:evaluation-scores). In most models, the AUROC scores peak in the middle layer. An exception is observed in Llama 3.1 8B Instruct, where evaluation scores (measured in AUROC, see 3.2) increase in the later layer.
(a) Comparison of PRISM (max) AUROC evaluation scores and PRISM polysemanticity scores across different models and layers.
To analyze the degree of polysemanticity, we use the gte-Qwen2-1.5B-instruct sentence transformer li-2023-GeneralTextEmbeddings to embed the five natural language descriptions generated per feature. We then compute pairwise cosine similarity within each set of five descriptions to estimate semantic consistency. In Figure LABEL:fig:poly-scores, polysemanticity scores show no consistent trend across layers and vary across models and feature types. Interestingly, we can also note that Gemma Scope SAE feature descriptions show the highest monosemanticity across all layers, as shown in Figure LABEL:fig:poly-scores. Despite high variability, our findings suggest that polysemanticity does not consistently increase or decrease with layer depth, and interpretability varies significantly across architectures.
@@ 6 Discussion and Conclusion
As we have focused on textual explanations, our descriptions are limited to concepts expressible in natural language, and thus may be unable to capture complex syntactic structures like graphs or algorithmic concepts where the underlying operation is not easily described in the constrained vocabulary space. Fixing the number of clusters in PRISM cannot guarantee that provided descriptions capture all or the most salient concepts for a feature; instead, they represent a subset of key concepts the feature is responsive to. Parametric measures such as MAD are sensitive to outliers, especially in NLP settings where activations often follow heavy-tailed distributions. To address this, we pair MAD with the more robust AUROC score, yielding a more comprehensive view of feature behavior. Finally, reliance on maximally activating corpus examples can restrict interpretations to observed concepts, limiting coverage of rare patterns or out-of-distribution effects across datasets bolukbasi_interpretability_2021.
@@
Capturing Polysemanticity with PRISM: A Multi-Concept Feature Description Framework